OpenStudy (anonymous):

base of a solid is enclosed by y=sqrt(x) y=2 and x=0 each cross-section is a semi-circle, how do I find the volume

OpenStudy (anonymous):

\[2\Pi \int\limits_{0}^{4} x dx\]

OpenStudy (anonymous):

this came from 2 * pi * r^2, the area of a circle. your radius in this case was sqrt(x), so you are technically integrating 2 * pi * sqrt(x)^2.

OpenStudy (kropot72):

Each cross-section is a semi-circle so the required integral would be:\[\pi \int\limits_{0}^{4}xdx\]

OpenStudy (anonymous):

 you need more info on this.. are the cross sections perpendicular to the x or y axis?